# Differentially Private SGD Without Clipping Bias: An Error-Feedback Approach

### v1

@@ Abstract
Differentially Private Stochastic Gradient Descent with gradient clipping (DPSGD-GC) is a powerful tool for training deep learning models using sensitive data, providing both a solid theoretical privacy guarantee and high efficiency. However, using DPSGD-GC to ensure Differential Privacy (DP) comes at the cost of model performance degradation due to DP noise injection and gradient clipping. Existing research has extensively analyzed the theoretical convergence of DPSGD-GC, and has shown that it only converges when using large clipping thresholds that are dependent on problem-specific parameters. Unfortunately, these parameters are often unknown in practice, making it hard to choose the optimal clipping threshold. Therefore, in practice, DPSGD-GC suffers from degraded performance due to the constant bias introduced by the clipping. In our work, we propose a new error-feedback (EF) DP algorithm as an alternative to DPSGD-GC, which not only offers a diminishing utility bound without inducing a constant clipping bias, but more importantly, it allows for an arbitrary choice of clipping threshold that is independent of the problem. We establish an algorithm-specific DP analysis for our proposed algorithm, providing privacy guarantees based on Rényi DP. Additionally, we demonstrate that under mild conditions, our algorithm can achieve nearly the same utility bound as DPSGD without gradient clipping. Our empirical results on Cifar-10/100 and E2E datasets show that the proposed algorithm achieves higher accuracies than DPSGD while maintaining the same level of DP guarantee.
@@ 1 Introduction
Challenges. In terms of theory, although the inclusion of clipping operation in DPSGD-GC ensures a strong DP guarantee, it considerably complicates the convergence analysis compared to the vanilla SGD algorithm. This is because the expected update direction, which is the expected clipped per-sample gradient in DPSGD-GC, may change dramatically, and it requires additional effort to analyze its alignment with the true gradient. Therefore, the early works on DPSGD with convergence analysis assume that the clipping threshold is chosen to be larger than the magnitude of each per-sample gradient, essentially making the clipping operation ineffective during training (Bassily et al., 2014; Wang et al., 2016; Feldman et al., 2020; Iyengar et al., 2019; Xu et al., 2021; Zhang et al., 2022; Li et al., 2022). Recent works use alternative assumptions and improve the convergence analysis for DPSGD-GC, but the convergence results still rely on an assumption-dependent choice of the clipping threshold (Fang et al., 2022; Chen et al., 2020; Yang et al., 2022; Qian et al., 2021; Zhang et al., 2020; Koloskova et al., 2023). However, the bounds in the assumptions of real-world problems are hard to estimate, and such a choice of clipping threshold is impossible to be satisfied in practice. Recent work Koloskova et al. (2023) has shown a negative result that, under the general assumptions for SGD, regardless of the choice of clipping threshold and stepsize, DPSGD-GC converges with a constant bias term, meaning in the limit the DPSGD-GC algorithm only converges to a neighborhood of the optimal or stationary solution. References Chen et al. (2020); Song et al. (2013) also provide a justification that the gradient clipping shifts the stationary solution of the original problem, thus causing an unavoidable constant bias (see our fixed-point analysis in Section 2.2).
@@
Our Contributions. In this work, we propose DiceSGD algorithm for DP training with both utility and privacy guarantees using a problem-independent clipping threshold. DiceSGD is motivated by the error-feedback (EF) mechanism – a classical procedure in signal processing (Howze & Bhattacharyya, 1997; Laakso & Hartimo, 1992) for cancelling quantization bias. Specifically, we propose a novel clipped EF mechanism which accumulates the error between the clipped update to the unclipped one at each iteration, and feeds the clipped error back to the next update. The proposed clipped EF mechanism satisfies the DP guarantee, while still preserving the ability to compensate for the per-sample gradient clipping bias and eventually eliminating the convergence bias caused by clipping. In contrast to existing works, the proposed DiceSGD provides DP guarantee and convergence guarantee without constant bias, while allowing a flexible choice of the clipping threshold. More importantly, we have observed that when the algorithm is applied to a number of applications, including image classification and natural language processing tasks, it does not suffer from performance degradation; nor does it require careful clipping threshold tuning, as compared with DPSGD.
We emphasize that the theoretical analysis for the proposed DiceSGD is challenging in the following sense: the clipping operation does not satisfy the firmly contracting assumption used in the typical analysis of EF algorithms; additionally, directly applying the conventional DP analysis to DiceSGD leads to an extremely loose privacy bound. Therefore, a new convergence and privacy analysis for the designed algorithm is required. We summarize our major contribution as follows:
@@
  • We provide the convergence proof for DiceSGD under either strong convexity or PL condition assumption, and show that DiceSGD eliminates the constant clipping bias compared with DPSGD-GC, when both using an arbitrary constant clipping threshold.
  • We develop an algorithm-specific Rényi-DP analysis for the proposed method, where the update consists of a privatized state and a non-privatized hidden state. We show that DiceSGD satisfies $(\epsilon,\delta)$-DP by injecting a slightly larger (i.e., a constant depending on the clipping threshold of the feedback error signal) DP noise compared with DPSGD-GC.
  • Finally, we perform rigorous empirical comparisons of our method to DPSGD-GC on a number of publicly available datasets, including Cifar-10/100 and E2E, to demonstrate the ability of our method to train models with a high privacy guarantee and good performance. Further, we conduct ablation studies on DiceSGD to show its insensitivity in the choice of hyper-parameters.
@@ 2 Preliminaries
In this section, we first formally introduce the formulation of the optimization problem and the definition of DP. Then we describe the assumptions, analyses, and issues of existing literature related to DPSGD and the EF mechanism.
@@ 2.1 Notations and assumptions
Symbol Meaning
$\mathbf{x},\mathbf{w},\mathbf{e}$ model variable, privacy noise, feedback error
$f(\cdot)$ objective function
$\mathcal{D},\mathcal{B},\xi$ dataset, minibatch, data sample
$N,B,i$ Dataset size, minibatch size, data index
$\mathbf{g}_{i},\nabla f(\mathbf{x};\xi_{i})$ gradient evaluated on $\xi_{i}$
$\epsilon,\delta$ parameters of $(\epsilon,\delta)$-DP
$T,t$ total iteration, iteration index
$\mu,L,\sigma$ strong convexity (PL condition), Lipschitz, variance constants
$C,C_{1},C_{2}$ clipping thresholds
Table 1: Symbols used in the main paper.
@@ 2.2 DPSGD-GC algorithm
Although providing a strong DP guarantee, the convergence property of DGSGD-GC is less satisfactory. Recent work Koloskova et al. (2023) has shown that without any extra assumption, DPSGD-GC with an arbitrary clipping threshold converges with a constant clipping bias, regardless of the convexity of the problem. Prior works that show the convergence of DPSGD-GC rely on extra assumptions on the problem and clipping thresholds that depend on these assumptions. Specifically, Chen et al. (2020) proves the convergence of DPSGD-GC under the assumption that the per-sample gradients have a symmetric distribution; Jin et al. (2022) give a high probability convergence result assuming that the per-sample gradients have a bounded domain and sufficiently large clipping threshold; work Yang et al. (2022) establishes the convergence of DPSGD-GC by assuming the deviation of per-sample gradient from the true gradient is bounded, and using a clipping threshold larger than the per-sample gradient deviation to ensure that clipped gradient “aligns” with the true gradient; light-tailed gradient variance assumption and a large clipping threshold has been used by Fang et al. (2022) to provide a high probability bound without constant bias.
@@
Although providing theoretically sound convergence analyses, the above-mentioned theoretical results do not provide practical guidance on choosing the clipping threshold in real-world applications. In these works, the choices of clipping thresholds depend on the problem parameters, which are hard or impossible to estimate. Therefore, these analyses cannot guarantee that clipping thresholds used in real-world training satisfy the requirements. Thus, DPSGD-GC still suffers from a constant clipping bias, and there is a strong need to design a new DP algorithm that does not suffer from clipping bias.
@@ 3.1 Algorithm description
Our DiceSGD algorithm is described in Algorithm 2. At round $t$, the algorithm first computes the update direction $\mathbf{v}^{t}$ by adding the clipped stochastic gradient with the clipped feedback error. Then, the algorithm updates the model parameters $\mathbf{x}^{t}$ with $\mathbf{v}^{t}$ and injects the DP noise $\mathbf{w}^{t}$. Finally, it computes the clipping error $\mathbf{e}^{t+1}$ for the next iteration. The algorithm only releases $\mathbf{x}^{t}$ at iteration $t$ and does not release $\mathbf{e}^{t}$ nor $\mathbf{v}^{t}$.
In the proposed algorithm, we introduce an extra variable $\mathbf{e}^{t}$ that records the clipping error. We keep it unclipped and privatize it when computing the update direction in the next iteration. Unlike the original EF mechanism, we do not feed it back directly to each per-sample gradient clipping operation (Line 5), as it breaks the sensitivity of the algorithm. Rather, we first clip $\mathbf{e}^{t}$ and add it to the averaged clipped gradient. Using such a clipped EF mechanism for privacy guarantee, we can balance the functionality of EF and the DP requirement of the algorithm.
@@
To see why the proposed algorithm has the potential to eliminate the clipping bias, let us again study the fixed-point of the DiceSGD algorithm. At the fixed-point, we have the following relation, where the expectation $\operatorname{\mathbb{E}}[\cdot]$ is taken on the randomness of the samples at the current iteration.
@@
which indicate that the fixed point of DiceSGD is given by
@@
We can show that when $C_{2}\geq C_{1},$ there exists $\mathbf{x},\mathbf{e}$ such that the fixed point is achieved. Specifically, $\mathbf{x}$ is chosen as $\nabla f(\mathbf{x})=0$; and $\mathbf{e}$ is chosen as $\mathbf{e}=-\frac{1}{N}\sum^{N}_{i=1}\mathrm{clip}\left(\nabla f(\mathbf{x};\xi_{i}),C_{1}\right).$ Because
@@
we have $\mathrm{clip}\left(\mathbf{e},C_{2}\right)=\mathbf{e},$ and the first equation is satisfied. These choices guarantee that the two equations are satisfied and the fixed point is achieved. The fixed-point analysis indicates that, unlike DPSGD-GC, as long as $C_{2}\geq C_{1}$ (which are problem independent), clipped EF can fully compensate the shift of the stationary solution caused by gradient clipping independent of any problem assumptions, and $\nabla f(\mathbf{x})=0$ is the fixed point of DiceSGD.
@@ 3.2 Theoretical analysis
In this section, we provide analysis for the proposed DiceSGD algorithm. We emphasize again that the challenge here is two-fold: 1) it is difficult to analyze convergence due to the combination of the EF mechanism and the clipping operation; 2) the DP analysis is non-trivial due to the presence of the non-privatized update of $\mathbf{e}^{t}$ as a hidden state. More specifically, the analyses of the convergence of the existing EF algorithms (Karimireddy et al., 2019; Li et al., 2022) relies on the assumption that the feedback error $\mathbf{e}^{t+1}$ in (3) is a firmly contractive mapping on $\mathbf{e}^{t}+\mathbf{g}^{t}$:
@@
Assumption 3.4 (PL Condition).
The gradient of the problem satisfies the following inequality:
$\frac{1}{2\mu}\left\lVert\nabla f(\mathbf{x})\right\rVert^{2}\geq f(\mathbf{x})-f^{\star},\forall~{}\mathbf{x}\in\mathbb{R}^{d}.$
The PL condition is also a widely used assumption when analyzing the convergence of SGD-type algorithms (Stich & Karimireddy, 2020; Lei et al., 2019; Karimi et al., 2016b). This assumption guarantees a quadratic growth to the gradient of the function. It is weaker than the strongly convex assumption as it only imposes a constraint based on the distance between one point to the set of optimal points.
@@
Assumptions 3.5 and 3.6 are commonly used for analyzing clipping operation (Zhang et al., 2020; Qian et al., 2021; Song et al., 2020), the convergence of DP algorithms (Yang et al., 2022), and distributed optimization (Li et al., 2022; Zhang et al., 2022). Assumption 3.5 assumes a smaller variance compared with the typical assumption (i.e., $\operatorname{\mathbb{E}}\left\lVert\nabla f(\mathbf{x})-\mathbf{g}_{i}\right\rVert^{2}\leq\sigma^{2}$), it implies that $\left\lVert\nabla f(\mathbf{x})-\mathbf{g}_{i}\right\rVert^{2}\leq\sigma^{2},\forall~{}i$, and it is necessary for bounding the clipping bias in the existing works (e.g.,in Yang et al. (2022)). Although these assumptions are also used in our analysis, contrasting with existing works, the clipping thresholds $C_{1},C_{2}$ in DiceSGD do not depend on $G$ or $\sigma$.
We now present the convergence theorem of the proposed DiceSGD algorithm under the strong convexity condition 3.3.
@@
Assume the problem satisfies Assumption 3.1, 3.2, 3.3, 3.5, and 3.6. Given any constant DP noise multiplier $\sigma_{1}$, by running DiceSGD (Algorithm 2) for $T$ iterations, choosing stepsize $\eta\leq\min\{\frac{1}{2\mu},\frac{3}{32L},\frac{\sqrt{C_{2}}}{5L\sqrt{C_{2}+T\max\{0,G^{\prime}-C_{2}\}}}\}$, $G^{\prime}=\max\{0,G+\sigma-C_{1}\}$, clipping thresholds $C_{2}\geq C_{1}>0$. It satisfies
$\operatorname{\mathbb{E}}[\mathcal{L}^{T+1}]\leq(1-\mu\eta)^{T}\mathcal{L}^{0}+\frac{\eta L}{2\mu}\left(d\sigma_{1}^{2}+\frac{5\sigma^{2}}{4B}+\frac{C_{1}^{2}}{2}\right)$ (4)
where we have defined
$\mathcal{L}^{t}:=f(\tilde{\mathbf{x}}^{t})-f^{\star}+\frac{\eta^{3}L^{2}}{6}\left\lVert\mathbf{e}^{t}\right\rVert^{2}+\frac{\eta(1-L\eta)}{2}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2},$
with $\tilde{\mathbf{x}}^{t}:=\mathbf{x}^{t}-\eta\mathbf{e}^{t}.$
@@
    We first define a virtual sequence $\{\tilde{\mathbf{x}}^{t}\},$ which is updated with the “unclipped gradients”, and we can apply the classic convergence analysis of SGD (without gradient clipping) for strongly convex problems on the virtual sequence. Due to the EF mechanism, the convergence result for DiceSGD has an extra term depending on $\operatorname{\mathbb{E}}\left\lVert\mathbf{e}^{t}\right\rVert^{2}$.
@@
    With the update of $\mathbf{e}^{t}$, we can derive a recursive bound on the extra term $\operatorname{\mathbb{E}}\left\lVert\mathbf{e}^{t}\right\rVert^{2}$ depending on $\left\lVert\mathbf{e}^{t-1}\right\rVert^{2}$ and the expected gradients. Unlike EF for contracting error which depends on the gradients with a constant factor independent of $T$, the error $\mathbf{e}^{t}$ caused by clipping operation gives an increasing factor of $T$ depending on the gradient.
@@
    By substituting the bound of $\mathbf{e}^{t}$ into the convergence result in step $1$, and choosing sufficiently small stepsize that compensates for the increasing dependency on $T$, we are able to derive a non-trivial convergence result for DiceSGD. A similar result can also be derived for problems satisfying Assumption 3.4.
Our analysis does not apply to the more general non-convex problem, because bounding the error term $\operatorname{\mathbb{E}}\left\lVert\mathbf{e}^{t}\right\rVert^{2}$ relies crucially on the strong convexity assumption. The detailed proof of the algorithm is given in Appendix A.1. A similar result also holds when replacing Assumption 3.3 with Assumption 3.4, and we provide the corresponding theorem and its proof in Appendix A.
Theorem 3.7 indicates that we can choose
$\eta=\min\left\{\frac{3}{32L},\frac{\sqrt{C_{2}}}{5L\sqrt{C_{2}+T\max\{0,G^{\prime}-C_{2}\}}}\},\frac{\log\left(2\mu^{2}T\mathcal{L}^{0}/L(d\sigma_{1}^{2}+\frac{5\sigma^{2}}{4B}+\frac{C_{1}^{2}}{2})\right)}{\mu T}\right\}.$
When $T$ is sufficiently large, $\eta=\mathcal{O}(\frac{\log(T)}{\mu T})$, so the overall convergence rate for DiceSGD is $\mathcal{O}\left(1-\frac{\log(T)}{T})^{T}\right)+\mathcal{O}\left(\frac{\log(T)}{T}\right)=\mathcal{O}\left(\frac{\log(T)}{T}\right)$.
Compared with the $\mathcal{O}(\frac{1}{T})$ lower bound convergence rate of DPSGD without gradient clipping under strong convexity (Bassily et al., 2014; Rakhlin et al., 2011), DiceSGD slows down by a factor of $\mathcal{O}(\log(T))$. However, compared with DPSGD-GC (Koloskova et al., 2023), DiceSGD fully eliminates the constant bias, and improves the convergence rate from $\mathcal{O}(1)$ to $\mathcal{O}(\frac{\log(T)}{T})$. The comparison is shown in Table 2.
@@
DPSGD $\mathcal{O}\left(\frac{G^{2}\log(1/\delta)}{N^{2}\epsilon^{2}}\right)$ $\mathcal{O}(\frac{G^{2}T\log(\frac{1}{\delta})}{N^{2}\epsilon^{2}})$ 3.3, 3.5, 3.6 $C\geq G+\sigma$
DPSGD-GC $\mathcal{O}\left(\frac{C^{2}\log(1/\delta)}{N^{2}\epsilon^{2}}\right)+\mathcal{O}(1)$ $\mathcal{O}(\frac{C^{2}T\log(\frac{1}{\delta})}{N^{2}\epsilon^{2}})$ 3.3, 3.5, 3.6 $C<G+\sigma$
DiceSGD $\mathcal{O}\left(\frac{\tilde{G}\log(N\epsilon)\log(1/\delta)}{N^{2}\epsilon^{2}}\right)$ $\mathcal{O}(\frac{\tilde{G}T\log(\frac{1}{\delta})}{N^{2}\epsilon^{2}})$ 3.3/3.4, 3.5, 3.6 Arbitrary
@@
Let us proceed with the privacy analysis of DiceSGD. We start with the notion of Rényi Differential Privacy (Mironov, 2017). By accounting for the distribution divergence of the stochastic gradient at iteration $t$ and the accumulated difference of $\mathbf{e}^{t}$ starting from $\mathbf{e}^{0}$, we are able to bound the Rényi divergence of $\mathbf{x}^{t+1}$ given two adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ and start with the same $\mathbf{x}^{t}$. Then by using the composition theorem of Rényi divergence, we provide the privacy guarantee for DiceSGD in the next result. Note that this result is independent of problem types as we do not assume either Assumption 3.3 and Assumption 3.4.

### v2

@@ Abstract
Differentially Private Stochastic Gradient Descent with Gradient Clipping (DPSGD-GC) is a powerful tool for training deep learning models using sensitive data, providing both a solid theoretical privacy guarantee and high efficiency. However, using DPSGD-GC to ensure Differential Privacy (DP) comes at the cost of model performance degradation due to DP noise injection and gradient clipping. Existing research has extensively analyzed the theoretical convergence of DPSGD-GC, and has shown that it only converges when using large clipping thresholds that are dependent on problem-specific parameters. Unfortunately, these parameters are often unknown in practice, making it hard to choose the optimal clipping threshold. Therefore, in practice, DPSGD-GC suffers from degraded performance due to the constant bias introduced by the clipping. In our work, we propose a new error-feedback (EF) DP algorithm as an alternative to DPSGD-GC, which not only offers a diminishing utility bound without inducing a constant clipping bias, but more importantly, it allows for an arbitrary choice of clipping threshold that is independent of the problem. We establish an algorithm-specific DP analysis for our proposed algorithm, providing privacy guarantees based on Rényi DP. Additionally, we demonstrate that under mild conditions, our algorithm can achieve nearly the same utility bound as DPSGD without gradient clipping. Our empirical results on standard datasets show that the proposed algorithm achieves higher accuracies than DPSGD while maintaining the same level of DP guarantee.
@@ 1 Introduction
Challenges. In terms of theory, although the inclusion of clipping operation in DPSGD-GC ensures a strong DP guarantee, it considerably complicates the convergence analysis compared to the vanilla SGD algorithm. This is because the expected update direction, which is the expected clipped per-sample gradient in DPSGD-GC, may change dramatically, and additional effort is required to analyze its alignment with the true gradient. Therefore, the early works on DPSGD with convergence analysis assume that the clipping threshold is chosen to be larger than the magnitude of each per-sample gradient, essentially making the clipping operation ineffective during training (Bassily et al., 2014; Wang et al., 2016; Feldman et al., 2020; Iyengar et al., 2019; Xu et al., 2021; Zhang et al., 2022; Li et al., 2022). Recent works use alternative assumptions and improve the convergence analysis for DPSGD-GC, but the convergence results still rely on an assumption-dependent choice of the clipping threshold (Fang et al., 2022; Chen et al., 2020; Yang et al., 2022; Qian et al., 2021; Zhang et al., 2020; Koloskova et al., 2023). However, the bounds in the assumptions of real-world problems are hard to estimate, and such a choice of clipping threshold is impossible to be satisfied in practice. Recent work Koloskova et al. (2023) has shown a negative result that, under the general assumptions for SGD, regardless of the choice of clipping threshold and stepsize, DPSGD-GC converges with a constant bias term, meaning in the limit the DPSGD-GC algorithm only converges to a neighborhood of the optimal or stationary solution. References Chen et al. (2020); Song et al. (2013) also provide a justification that the gradient clipping shifts the stationary solution of the original problem, thus causing an unavoidable constant bias (see our fixed-point analysis in Section 2.2).
@@
Our Contributions. In this work, we propose DiceSGD algorithm for DP training with both utility and privacy guarantees using a problem-independent clipping threshold. DiceSGD is motivated by the error-feedback (EF) mechanism – a classical procedure in signal processing (Howze & Bhattacharyya, 1997; Laakso & Hartimo, 1992) for cancelling quantization bias. Specifically, we propose a novel clipped EF mechanism which accumulates the error between the clipped update to the unclipped one at each iteration, and feeds the clipped error back to the next update. The proposed clipped EF mechanism satisfies the DP guarantee, while still preserving the ability to compensate for the per-sample gradient clipping bias and eventually eliminating the convergence bias caused by clipping. In contrast to existing works, the proposed DiceSGD provides DP guarantee and convergence guarantee without constant bias, while allowing a flexible choice of the clipping threshold. More importantly, we have observed that when the algorithm is applied to a number of applications, including image classification and natural language processing tasks, it does not suffer from performance degradation; nor does it require careful clipping threshold tuning.
We emphasize that the theoretical analysis for the proposed DiceSGD is challenging in the following sense: the clipping operation does not satisfy the firmly contracting assumption used in the typical analysis of EF algorithms; additionally, directly applying the conventional DP analysis to DiceSGD leads to an extremely loose bound. Therefore, a new convergence and privacy analysis for the designed algorithm is required. We summarize our major contribution as follows:
@@
  • We provide the convergence proof for DiceSGD under general non-convex and Lipschitz-smooth assumption, and show that DiceSGD eliminates the constant clipping bias compared with DPSGD-GC with an arbitrary constant clipping threshold.
  • We develop an algorithm-specific Rényi-DP analysis for the proposed method, where the update consists of a privatized state and a non-privatized hidden state. We show that DiceSGD satisfies $(\epsilon,\delta)$-DP by injecting a slightly (i.e., a constant depending on the clipping threshold of the feedback error signal) larger DP noise compared with DPSGD-GC.
  • Finally, we perform rigorous empirical comparisons of our method to DPSGD-GC on a number of publicly available datasets to demonstrate the ability of our method to train models with a high privacy guarantee and good performance. Further, we conduct ablation studies on DiceSGD to show its stability in the choice of hyper-parameters.
@@ 2.2 DPSGD-GC algorithm
Although providing a strong DP guarantee, the convergence property of DGSGD-GC is less satisfactory. Recent work Koloskova et al. (2023) has shown that without any extra assumption, DPSGD-GC with an arbitrary clipping threshold converges with a constant clipping bias, regardless of the convexity of the problem. Prior works that show the convergence of DPSGD-GC rely on extra assumptions on the problem and clipping thresholds that depend on these assumptions. Specifically, Chen et al. (2020) proves the convergence of DPSGD-GC under the assumption that the per-sample gradients have a symmetric distribution; Jin et al. (2022) gives a high probability convergence result assuming that the per-sample gradients have a bounded domain and sufficiently large clipping threshold; Yang et al. (2022) establishes the convergence of DPSGD-GC by assuming that the deviation of per-sample gradient from the true gradient is bounded, and using a clipping threshold larger than the per-sample gradient deviation to ensure that clipped gradient “aligns” with the true gradient; light-tailed gradient variance assumption and a large clipping threshold has been used by Fang et al. (2022) to provide a high probability bound without constant bias.
@@
Although providing theoretically sound convergence analyses, the theoretical results in Chen et al. (2020); Jin et al. (2022); Yang et al. (2022); Fang et al. (2022) do not provide practical guidance on choosing the clipping threshold in real-world applications. In these works, the choices of clipping thresholds depend on the problem parameters, which are hard or impossible to estimate. Therefore, these analyses cannot guarantee that clipping thresholds used in real-world training satisfy the requirements. Thus, DPSGD-GC still suffers from a constant clipping bias, and there is a strong need to design a new DP algorithm that does not suffer from clipping bias.
@@ 3.1 Algorithm description
Our DiceSGD algorithm is described in Algorithm 2 and Figure 1. At round $t$, the algorithm first computes the update direction $\mathbf{v}^{t}$ by adding the clipped stochastic gradient with the clipped feedback error. Then, the algorithm updates the model parameters $\mathbf{x}^{t}$ with $\mathbf{v}^{t}$ and injects the DP noise $\mathbf{w}^{t}$. Finally, it computes the clipping error $\mathbf{e}^{t+1}$ for the next iteration. The algorithm only releases $\mathbf{x}^{t}$ at iteration $t$ and does not release $\mathbf{e}^{t}$ nor $\mathbf{v}^{t}$.
Figure 1: The flow diagram of DiceSGD. The clipped EF components are highlighted in red, and DP components are marked in yellow. $z^{-1}$ denotes the unit delay.
In the proposed algorithm, we introduce an extra variable $\mathbf{e}^{t}$ that records the clipping error. We keep it unclipped and privatize it when computing the update direction in the next iteration. As an important algorithm design consideration for DP requirement, unlike the original EF mechanism, we do not feed $\mathbf{e}^{t}$ back directly to each per-sample gradient clipping operation (Line 5), because it would break the sensitivity of the algorithm. Rather, we first clip $\mathbf{e}^{t}$ and add it to the averaged clipped gradient. Using such a clipped EF mechanism for privacy guarantee, we can balance the functionality of EF and the DP requirement of the algorithm.
@@
To see why the proposed algorithm has the potential of eliminating the clipping bias, let us again study the fixed-point of the DiceSGD algorithm. At the fixed-point, we have the following relation, where the expectation $\operatorname{\mathbb{E}}[\cdot]$ is taken on the randomness of the samples at the current iteration.
@@
which indicates that the fixed-point of DiceSGD is given by
@@
We can show that when $C_{2}\geq C_{1},$ there exists $\mathbf{x},\mathbf{e}$ such that the fixed point is achieved. Specifically, by choosing $\mathbf{x}$ satisfies $\nabla f(\mathbf{x})=0$; and $\mathbf{e}$ is chosen as $\mathbf{e}=-\frac{1}{N}\sum^{N}_{i=1}\mathrm{clip}\left(\nabla f(\mathbf{x};\xi_{i}),C_{1}\right).$ Note that
@@
we have $\mathrm{clip}\left(\mathbf{e},C_{2}\right)=\mathbf{e}$ as long as $C_{2}\geq C_{1}$, and the first equation is satisfied. These choices guarantee that the two equations are satisfied and the fixed point is achieved. The fixed-point analysis indicates that, unlike DPSGD-GC, as long as $C_{2}\geq C_{1}$, a condition that is problem independent, clipped EF can potentially fully compensate the shift of the stationary solution caused by gradient clipping independent of any problem assumptions, and $\nabla f(\mathbf{x})=0$ is the fixed point of DiceSGD.
@@ 3.2 Theoretical analysis
In this section, we provide analysis for the proposed DiceSGD algorithm. We emphasize again that the challenge here is two-fold: 1) it is difficult to analyze convergence due to the combination of the EF mechanism and the clipping operation; 2) the DP analysis is non-trivial due to the presence of the non-privatized update of $\mathbf{e}^{t}$ as a hidden state. To see the first challenge, more specifically, the analyses of the convergence of the existing EF algorithms (Karimireddy et al., 2019; Li et al., 2022) relies on the assumption that the feedback error $\mathbf{e}^{t+1}$ in (3) is a firmly contractive mapping on $\mathbf{e}^{t}+\mathbf{g}^{t}$:
@@
Assumptions 3.4 and 3.5 are commonly used for analyzing clipping operation (Zhang et al., 2020; Qian et al., 2021; Song et al., 2020), the convergence of DP algorithms (Yang et al., 2022), and distributed optimization (Li et al., 2022; Zhang et al., 2022). Assumption 3.4 assumes a smaller variance compared with the typical assumption (i.e., $\operatorname{\mathbb{E}}\left\lVert\nabla f(\mathbf{x})-\mathbf{g}_{i}\right\rVert^{2}\leq\sigma^{2}$), it implies that $\left\lVert\nabla f(\mathbf{x})-\mathbf{g}_{i}\right\rVert^{2}\leq\sigma^{2},\forall~{}i$, and it is necessary for bounding the clipping bias in the existing works (e.g.,in Yang et al. (2022)). Although these assumptions are also used in our analysis, contrasting with existing works, the clipping thresholds $C_{1},C_{2}$ in DiceSGD do not depend on $G$ or $\sigma$.
We now present the convergence theorem of the proposed DiceSGD algorithm under the non-convex smooth setting Assumption 3.2:
@@
Assume the problem satisfies Assumption 3.1, 3.2, 3.4, and 3.5. Given any constant DP noise multiplier $\sigma_{1}$, by running DiceSGD (Algorithm 2) for $T$ iterations, choosing stepsize $\eta=\sqrt{\frac{2(f(\mathbf{x}^{0})-f^{\star})}{TL(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2})}}$, clipping thresholds $C_{2}\geq 3C_{1}+\frac{\sigma}{B}>0$. It satisfies
$\operatorname{\mathbb{E}}_{t}\left[\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}\right]$ $\leq 2\sqrt{\frac{2L(f(\mathbf{x}^{0})-f^{\star})(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2})}{T}},$ (4)
where the expectation $\operatorname{\mathbb{E}}_{t}$ is taken over $t\in\{0,\dots,T-1\}$, following distribution $\frac{A_{t}}{\sum^{T-1}_{t=0}A_{t}}$, with $\{A_{t}\}\in(0,1]$ being a strictly positive sequence defined in (12), Appendix A.
@@
    We first apply the convergence analysis of biased SGD for non-convex problems with update direction $\operatorname{\mathbb{E}}[\mathbf{v}^{t}]$. Due to the EF mechanism, the convergence result for DiceSGD directly depends on the recursion of $\mathbf{e}^{t}$, which corrects the bias at iteration $t-1$.
@@
    With the update of $\mathbf{e}^{t}$, we can derive a recursive bound on the key term $\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}[\mathbf{e}^{t}]\right\rangle$. Unlike EF for contracting error, which depends on the gradients with a constant factor independent of $T$, the error $\mathbf{e}^{t}$ caused by clipping operation requires a much tighter recursion directly on the inner product between $\mathbf{e}^{t}$ and $\nabla f(\mathbf{x}^{t})$ for analysis. And the coefficients before the gradient heavily depend on the clipping factor.
@@
    By substituting the bound of $\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}[\mathbf{e}^{t}]\right\rangle$ into the convergence result in step $1$, and choosing sufficiently small stepsize and adequate clipping factor ratio that compensates for the stochastic noise and the clipping bias, we are able to derive a non-trivial convergence result for DiceSGD.
Theorem 3.6 indicates that the overall convergence rate for DiceSGD is $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$ for the general non-convex setting, which matches the $\mathcal{O}(\frac{1}{\sqrt{T}})$ lower bound convergence rate of DPSGD without gradient clipping under non-convexity (Bassily et al., 2014; Rakhlin et al., 2011). However, compared with DPSGD-GC (Koloskova et al., 2023), DiceSGD fully eliminates the constant bias and improves the convergence rate from $\mathcal{O}(1)$ to $\mathcal{O}(\frac{1}{\sqrt{T}})$. The comparison is shown in Table 1.
@@
DPSGD $\mathcal{O}\left(\frac{G\sqrt{\log(1/\delta)}}{N\epsilon}\right)$ $\mathcal{O}(\frac{G\sqrt{T\log(\frac{1}{\delta})}}{N\epsilon})$ 3.4, 3.5 $C\geq G+\sigma$
DPSGD-GC $\mathcal{O}\left(\frac{C\sqrt{\log(1/\delta)}}{N\epsilon}\right)+\mathcal{O}(1)$ $\mathcal{O}(\frac{C\sqrt{T\log(\frac{1}{\delta})}}{N\epsilon})$ 3.4, 3.5 $C<G+\sigma$
DiceSGD $\mathcal{O}\left(\frac{\sqrt{\tilde{G}\log(1/\delta)}}{N\epsilon}\right)$ $\mathcal{O}(\frac{\sqrt{\tilde{G}T\log(\frac{1}{\delta})}}{N\epsilon})$ 3.4, 3.5 Independent of $G$
@@
Let us proceed with the privacy analysis of DiceSGD. We start with the notion of Rényi Differential Privacy (Mironov, 2017). By accounting for the distribution divergence of the stochastic gradient at iteration $t$ and the accumulated difference of $\mathbf{e}^{t}$ starting from $\mathbf{e}^{0}$, we are able to bound the Rényi divergence of $\mathbf{x}^{t+1}$ given two adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ and start with the same $\mathbf{x}^{t}$. Then by using the composition theorem of Rényi divergence, we provide the privacy guarantee for DiceSGD in the next result.